Bill and Mary filed a joint Federal income tax return this year. Mary owns a 30% interest in MAJIC Partnership, a women's dress boutique. Mary's share of the partnership's net income is $280,000. Her shares of the partnership's W–2 wages and unadjusted basis of depreciable property are $100,000 and $300,000, respectively.

If an amount is zero, enter "0".

a. What is Bill and Mary's maximum QBI deduction if their total taxable income is $300,000?

b. What is the maximum QBI deduction if Bill and Mary's total taxable income is $450,000?
Answer #1

a) If taxable income is less than $315,000 as a married filing jointly tax payer then the person can take straight 20% deduction

Therefore 20% of QBI = $280000 * 20% = $56,000 is the maximum deduction

b) If taxable income is more than $315,000 as a married filing jointly tax payer limitation to be calculated

QBI deduction is limited to the least of

i) 20% of QBI or

ii) 50% of the company's W-2 wages or sum of 25% of W-2 wages and 2.5% of the unadjusted basis of all qualified property whichever is more beneficial to the tax payer

Therefore least of

i) 20% of QBI = $280,000 * 20% = $56,000

ii) 50% of W-2 wages = $100,000 * 50% = $50,000 or

25% of W-2 wages plus 2.5% of unadjusted basis of depreciable property = (25% * 100,000)+(2.5% * 300,000)

=$25,000+$7,500

=$32,500

Therefore QBI maximum deduction is $ 50,000